sin 70 degrees into sin 20 degree minus cos 20 degree into cosec 70 degree

Grade:10

2 Answers

Vikas TU
6 years ago
.sec 70.sin 20 – cos 20. Cosec70 
=1/cos 70 *sin(90-70)- cos(90-70).1/sin 70 
=1/cos 70 * cos 70-sin 70 * 1/sin 70 
=1-1 
=0 
 
manish raj
6 years ago
tan20degree+2tan50degree=====================================================================================================================================================================================================================================================================